OpenStudy (anonymous):

I really really need help with this question, so I am reposting it. Solve the equation (3x-1)/(x+7) -(1)/(3x+2) = (5-x)/(3x^2+23x+14)

OpenStudy (inkyvoyd):

You had a lot of great answers on your first post. Check back on them.

OpenStudy (anonymous):

well I typically know how to do these problems...however I keep getting a different answer. I was hoping for opinions.

OpenStudy (anonymous):

When I first saw the problem, I thought that I should clear fractions with a common LCD The common LCD is (3x+2)(x+7) so multiply everything by that. then I got (3x-1)(3x+2)-(x+7)=(5-x) when I foil and combine like terms, setting the entire thing equal to 0, I got: (3x+5)(3x-2)=0 so x= 2/3 and -5/3 That is what I have...not sure if it is correct tho..

OpenStudy (dumbcow):

this looks good --> (3x-1)(3x+2)-(x+7)=(5-x) something wrong happened here though (3x+5)(3x-2)=0 (3x-1)(3x+2) = 9x^2 +3x -2 (5-x) +(x+7) = 12 --> 9x^2 +3x -14 = 0

OpenStudy (anonymous):

how do I factor something like that? :O

OpenStudy (dumbcow):

in this case it doesn't factor, you have to use quadratic formula

OpenStudy (anonymous):

ah...okay :) I remember how to use that THANKS SO MUCH!!!!! you are my new fan <3 <3 <3

OpenStudy (dumbcow):

:)